FYI we plan to switch the suspicious 1471N module that holds the TPC driftvoltages. regards Flemming ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov ----- Original Message ----- From: "Angelika Drees" <drees@bnl.gov> Sent: Thursday, January 10, 2002 4:03 PM Subject: Controlled Access tomorrow > Hi everybody, > > (be prepared to get multiple emails from me every day if we get even > closer to the end of this run ;) ) > ramp tuning last night was pretty successful :). We had 2 good stores > with more than 3 Tp per ring. At the end of the last store the yellow > polarimeter target broke. By now we switched to target #3 in yellow. > For the moment we're at injection and plan to study beam losses in IR10 > as soon as we get good injection and have the next ramp. For the rest > of today and the night we should be in physics running mode. > > Tomorrow morning we'll have a CONTROLLED access in IR6 and sector 2 for > 1 h(!). This access will be granted pending a store dump but should > happen between 8:00 and 12:00. In order to minimize the time it takes us > to come back, no other area should be opened. So far I havn't heard from > any other need though. > > Have a nice afternoon, > > Angelika > > > >
This archive was generated by hypermail 2b30 : Thu Jan 10 2002 - 21:48:20 EST